OpenStudy (yellowlegoguy99):

Which operation would you use on the exponents to simplify the expression 5^7 × 5^11? Addition Subtraction Multiplication Division

OpenStudy (yellowlegoguy99):

@SolomonZelman @shiraz14

OpenStudy (amistre64):

if we have 7, xs, multiplied together, and also have 11, xs, multiplied together; then what is the total number of xs that are multiplied together ?

OpenStudy (yellowlegoguy99):

So it's C?

OpenStudy (amistre64):

.... how many xs do we have multiplied together? how do we determine the number of xs entotal that are multiplied together?

OpenStudy (yellowlegoguy99):

xs....?

sammixboo (sammixboo):

So do you know PEMDAS and The Law of Exponents?

OpenStudy (amistre64):

yeah, x^7 means: x*x*x*x*x*x*x ... 7 xs multiplied togehter

OpenStudy (amistre64):

if we have 7 xs multiplied together, and 11 xs multiplied together, then how many xs entotal are mutliplied togther, and how do we figure that total out?

OpenStudy (yellowlegoguy99):

Duh I know pemdas and the law of exponents

OpenStudy (yellowlegoguy99):

(5 x 5 x 5 x 5 x 5 x 5 x 5) x (5 x 5 x 5 x 5 x 5 x 5 x 5 x 5 x 5 x 5 x 5)

OpenStudy (yellowlegoguy99):

78125 x 48828125

sammixboo (sammixboo):

Ok now how many 5's are there in that?

OpenStudy (yellowlegoguy99):

7 5's and 11 5's

OpenStudy (amistre64):

what operation can we use to determine the total number of 5s now?

OpenStudy (the_fizicx99):

I don't see how PEMDAS has anything to do with this, it's follows an exponent law. \(\ \sf 5^7 = 5 \times 5 \times 5\times 5\times 5\times 5 \times 5\) \(\ \sf 5^{11} = 5 \times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\) So, \(\ \sf 5 \times 5 \times 5\times 5\times 5\times 5 \times 5\) \(\ \sf 5 \times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\times 5\) = 5^? Or, simply apply the law of exponents, \(\ \sf \Large x^2 \times x^3 = x^{2+3} = x^5 \)

sammixboo (sammixboo):

See this is what you do: 5^7 x 5^11 - (5x5x5x5x5x5x5) x (5x5x5x5x5x5x5x5x5x5x5) Count how many fives are there ALL TOGETHER. There is 18 5's What did you do to know how many 5's where there? Basically you are simplifying the exponents into one exponent.

OpenStudy (yellowlegoguy99):

5^18

sammixboo (sammixboo):

I know @tHe_FiZiCx99 I just read the problem wrong at first.. Now @yellowlegoguy99 what did you do to get 5^18?

OpenStudy (the_fizicx99):

Yes, \(\ \sf 5^7 \times 5^{11} = 5^{7 + 11} = 5^{18} \)

sammixboo (sammixboo):

@yellowlegoguy99 What did you do to get 5^18... You added/multiplied/divided/subtracted (choose which one) 7 and 11 to get 18